A cars bumper is designed to withstand a 6.84-km/h (1.9-m/s) collision with avable object without damage to the body of the car. The bumper cushions the shock by absorbing the force over a distance. Calculate the magnitude of the average force on a bumper that collapses 0.165 m while bringing a 950-kg car to rest from an initial speed of 1.9 m/s
Answer #1

Solution)

We know,

Acceleration, a = (v^2/2d) = (1.9^2/2*0.165) = 10.93 m/sec^2

Now, we know

Force, F = ma = 950 x 10.93 = 10392.42 N (Ans)

=======

Good luck!:)